OpenStudy (anonymous):

What is the distance between P(-3, -2) and Q(5, -2)? 12 14 10 8 2. What is the distance between P(7, 0) and Q(8, 0)? 1 4 2 3 3. What is the distance between points P(0, 4) and Q(3, 0)? 8 5 6 7 4. What is the distance between points P(1, 7) and Q(9, -8)? 15 13 17 11 5. Find the midpoint of segment P Q if P(4, 1) and Q(12, -5)? (2, -8) (8, -2) (-8, -2) (-2, 8) 6. Find the midpoint of segment P Q if P(10, 2) and Q(8, 0)? (9, 1) (-9, -1) (1, 9) (-1, 9) 7.

OpenStudy (anonymous):

Please teach me how to do this!

OpenStudy (johnweldon1993):

We can use the distance formula \[\large \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\]

OpenStudy (anonymous):

So the equation will be (-3-2)^2 + ( -2+2)^2 ? @ohnweldon1993

OpenStudy (johnweldon1993):

For example with #1 "What is the distance between P(-3, -2) and Q(5, -2)?" \(\large P(\color \red{-3}, \color \purple{-2}) Q(\color \green{5},\color \yellow{-2})\) \[\large (\color \red{x_1}, \color \purple{y_1}) Q(\color \green{x_2},\color \yellow{y_2})\] so we would have \[\large \sqrt{(5 - (-3))^2 + (-2 - (-2))^2}\]

OpenStudy (johnweldon1993):

Does that make sense?

OpenStudy (johnweldon1993):

So in simplifying, we have \[\large \sqrt{(5 + 3)^2 + (-2 + 2)^2}\] \[\large \sqrt{(8)^2 + (0)^2}\] \[\large \sqrt{64} = 8\]

OpenStudy (johnweldon1993):

And then for the midpoints on the last 2 questions we can basically do the same thing \[\large Midpoint = (\frac{x_1 + x_2}{2} , \frac{y_1 + y_2}{2})\]

OpenStudy (anonymous):

yes, I understand the first one! thank you. And the midpoint one... will you show me how to do it? @johnweldon1993

OpenStudy (johnweldon1993):

Of course :) so lets use the last problem as the walkthrough "Find the midpoint of segment P Q if P(10, 2) and Q(8, 0)?" So we'll set it up again \[\large P(\color \red{10}, \color \purple{2}) Q(\color \green{8},\color \yellow{0})\] \[\large P(\color \red{x_1}, \color \purple{y_1}) Q(\color \green{x_1},\color \yellow{y_2})\] so for the mid point we use the formula I wrote up there \[\large Midpoint = (\frac{x_1 + x_2}{2} , \frac{y_1 + y_2}{2})\] \[\large Midpoint = (\frac{10 + 8}{2} , \frac{2 + 0}{2})\] \[\large Midpoint = (\frac{18}{2} , \frac{2}{2})\] \[\large Midpoint = (9 , 1)\]

OpenStudy (johnweldon1993):

Sorry, up there is a little typo it should be \[\large P(\color \red{x_1}, \color \purple{y_1}) Q(\color \green{x_2},\color \yellow{y_2})\] *Just accidentally wrote \(\large x_1\) twice*

OpenStudy (anonymous):

I got it :) Thank you so much!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! @johnweldon1993

OpenStudy (johnweldon1993):

Anytime hun :)

OpenStudy (anonymous):

did you receive a medal? @johnweldon1993

OpenStudy (johnweldon1993):

Oh, yes I did @chappy0729 thanks for that, dont do it for the medal but it was very nice of you ^_^

OpenStudy (anonymous):

hahaha okay:) well thank you for helping me ! @johnweldon1993

OpenStudy (johnweldon1993):

Of course :) if you need anything else feel free to message me :)

OpenStudy (anonymous):

Okay :) I will:p @johnweldon1993

OpenStudy (johnweldon1993):

:P